### v1

# Hybrid Near-field and Far-field Localization with Holographic MIMO

Abstract
Due to its ability to precisely control wireless beams, holographic multiple-input multiple-output (HMIMO) is expected to be a promising solution to achieve high-accuracy localization. However, as the scale of HMIMO increases to improve beam control capability, the corresponding near-field (NF) region expands, indicating that users may exist in both NF and far-field (FF) regions with different electromagnetic transmission characteristics. As a result, existing methods for pure NF or FF localization are no longer applicable. We consider a hybrid NF and FF localization scenario in this paper, where a base station (BS) locates multiple users in both NF and FF regions with the aid of a reconfigurable intelligent surface (RIS), which is a low-cost implementation of HMIMO. In such a scenario, it is difficult to locate the users and optimize the RIS phase shifts because whether the location of the user is in the NF or FF region is unknown, and the channels of different users are coupled. To tackle this challenge, we propose a RIS-enabled localization method that searches the users in both NF and FF regions and tackles the coupling issue by jointly estimating all user locations. We derive the localization error bound by considering the channel coupling and propose an RIS phase shift optimization algorithm that minimizes the derived bound. Simulations show the effectiveness of the proposed method and demonstrate the performance gain compared to pure NF and FF techniques.
Index Terms:
Holographic multiple-input multiple-output (HMIMO), reconfigurable intelligent surface (RIS), near-field localization, far-field localization.
I Introduction
Future sixth-generation (6G) wireless systems will support a wide range of applications such as autonomous driving and augmented reality[2] that require not only ubiquitous communication but also user localization capabilities. Consequently, research interest in advanced localization methods has surged, among which holographic multiple-input multiple-output (HMIMO)-based localization methods are expected to play a critical role due to its potential to achieve high localization accuracy. Specifically, HMIMO is an electromagnetic surface containing a massive number of antennas or reconfigurable metamaterial elements[3][4]. Due to the large number of tunable elements in the surface, HMIMO can precisely manipulate wireless beams to accurately locate users.
Traditional HMIMO-based localization works mainly fall into two categories: far-field (FF)[5, 6, 7] and near-field (NF) localization schemes[8, 9, 10]. In the FF region, the distances between the HMIMO and users are sufficiently large, and the signal channel is effectively described using a plane wave model. In [5], the authors considered FF user localization aided by millimeter wave (mmWave) HMIMO, where the users receive the signal transmitted by the BS to locate themselves. A maximum likelihood (ML) location estimator and two sub-optimal estimators with lower complexity were derived. HMIMO-based localization without base stations (BSs) was also investigated in [6], where the user transmitted the signal and received the signal reflected from the HMIMO. In [11], the authors considered the scenario that both the LOS path and the reflected path via the HMIMO exist simultaneously and proposed a low-complexity ML estimator.
HMIMO-based NF localization schemes have attracted much attention recently due to an increasing interest in large-scale HMIMO. Specifically, to improve beam control capability, the scale of HMIMO is enlarged, leading to the expansion of the NF region[12]. Different from the FF case, the spherical wave model has to be employed to characterize the NF channel because the signal transmission characteristics of the NF are determined by both the range and angle of the user location. In [10], the authors considered the scenario where the LOS path between the BS and the user was blocked and used HMIMO to construct virtual line-of-sight (VLOS) paths between the users and the BS. A localization scheme targeting NF wireless users was investigated. In [9], a HMIMO functioning as a lens receiver was utilized to achieve NF localization. The authors in [8] developed a received signal strength (RSS)-based localization method for NF users and a RIS phase shift adjustment method to enhance accuracy.
In practical scenarios, users may exist in both the NF and FF regions of the HMIMO. Existing algorithms for pure NF or FF users are therefore not always applicable. Specifically, FF localization methods suffer from low localization accuracy for NF user localization due to the significant error in describing NF channels using the plane wave model. In contrast, NF localization techniques are not ideal for FF users because the coupling of range and angle in the spherical model brings extra complexity, which has two main effects: first, it prolongs the running time of the algorithm; second, the more complex NF model will hinder convergence of the method and lead to accuracy degradation. In addition, the localization gain for the FF users brought by the NF model is limited. Hence, a localization scheme is required to adapt to the localization of both the NF and FF users.
In this paper, a HMIMO-based localization scheme is proposed for users in both NF and FF regions. Specifically, in the considered scenario, a base station (BS) and a reconfigurable intelligent surface (RIS) cooperate to locate multiple users in a multipath environment. The RIS is widely acknowledged as a low-cost implementation of HMIMO that can create a customized reflection beam by varying the phase shifts of its reflection elements. In the localization process, the RIS reflects the signals emitted from several single-antenna users to the BS. Then, the BS uses the received signal to estimate the locations of the users in both NF and FF regions. The phase shifts of the RIS are optimized based on the estimated locations of the users. The location estimation and RIS phase shift optimization are performed iteratively to improve localization accuracy.
Several challenges in this scenario need to be addressed. First, it is unknown whether each user or scatter is in the NF or the FF region, which leads to difficulty in designing the localization algorithm. Second, the channels of different users are coupled due to the existence of common scatters and shared RIS phase shifts. To tackle the above challenges efficiently, a hybrid NF and FF localization algorithm is proposed, in which we first sample the NF and FF regions separately, and then compare the received signals with the reconstructed signals at the sampled locations to locate the users and scatters. The locations of the users and scatters are jointly estimated by considering the channel coupling effect. Then, a RIS phase shift optimization problem is formulated to minimize the sum of the Cramér-Rao bounds (CRBs) of the user localization errors. A complex circle manifold-based method is proposed to solve the optimization problem with the coupled CRBs.
Our contributions can be summarized as follows.
  • We consider a hybrid NF and FF localization scenario, where the BS locates multiple users in hybrid regions with the aid of an RIS. A hybrid NF and FF localization framework is proposed, where the localization of the users and the RIS phase shifts optimization are iteratively conducted to improve localization accuracy.
  • Based on the principle of minimizing localization loss, the location estimation problem is formulated and a multi-user localization algorithm is developed, which localizes users in the hybrid NF and FF region under the channel coupling effect.
  • We formulate the RIS phase shift optimization problem to minimize the sum of CRBs of the location estimation errors. A complex circle manifold-based algorithm is then devised to obtain optimal RIS phase shifts.
  • We compare the performance of the proposed approach and other RIS-enabled localization techniques. Simulation results show that the proposed approach outperforms other methods by more than 30% in the root mean square error (RMSE) of angle estimation for both NF and FF users and by 60% in the RMSE of range estimation for NF users given $-10\leq\text{SNR}\leq 0$.
The rest of this paper is organized as follows. The localization scenario, the signal model, and the localization protocol are provided in Sec. II. In Sec. III, we propose the localization algorithm. Then, the RIS phase shift optimization algorithm is designed in Sec. IV. The performance of the proposed method is analyzed in Sec. V. In Sec. I, we present the simulation results and discussions. Finally, conclusions are drawn in Sec. VII.
Notation: The transpose and conjugate transpose are denoted by $(\cdot)^{T}$ and $(\cdot)^{H}$, respectively. We use $\|\bm{x}\|_{2}$ and $\|\bm{x}\|_{0}$ mean the $l_{2}$ and $l_{0}$ norm of vector $\bm{x}$, and ||$\cdot$|| ⋅ | denotes the number of elements in a set. Unbolded $x$, bold lower-case $\bm{x}$ and bold upper-case $\bm{X}$ denote scalar, vector and matrix, respectively. The collection of all complex numbers is denoted by $\mathbb{C}$. The Hadamard product is represented by $\odot$.
II System Model
In this section, we first introduce the scenario of multi-user localization. Then, we describe the signal model and introduce a localization protocol for the proposed scenario.
II-A Localization Scenario
As shown in Fig. 1, we consider a localization scenario consisting of $K$ users, a RIS, and a BS. Users and the BS are each equipped with a single antenna. The RIS containing $N=N_{1}\times N_{2}$ elements is placed on the Y-Z plane, where the center of the RIS is at the origin of the coordinate system. There also exist obstacles, which reflect the signals and create multipath channels. We assume the line-of-sight (LOS) path between each user and the BS is blocked, and there only exist reflection paths via the RIS between the users and the BS.
During the localization process, each user sends a narrowband signal to the RIS, and the RIS reflects it to the BS. To avoid multi-user interference, frequency division multiplexing (FDM) is applied to transmit signals from different users, i.e. the $k$-th user sends the signal over frequency $f_{k}$, and we have $f_{k}\neq f_{k}^{\prime}$ ($k\neq k^{\prime}$). The BS then estimates the locations of the users using the received signals. We assume the locations of the RIS and the BS are known, while the locations of the users and scattering centers in the obstacles [13] are unknown. Whether the users are located in the NF or the FF regions of the RIS is also unknown.
Figure 1: RIS-enabled multi-user localization system.
II-B Signal Model
In the uplink transmission of the $k$-th user, the signal $y_{k}$ received by the BS is given by
$y_{k}=\bm{\beta}^{T}\bm{h}_{k}(\bm{p}_{k},\bm{Q})s_{k}+\epsilon,$ (1)
where $\bm{\beta}=[\beta_{1},\beta_{2},...,\beta_{N}]^{T}\in\mathbb{C}^{N\times 1}$ is the phase shift vector of the RIS which is shared for all $K$ users, $\beta_{n}$ represents the phase shift of the $n$-th RIS element with $|\beta_{n}|=1$, $s_{k}\in\mathbb{C}^{1\times 1}$ represents the narrowband signal transmitted by the $k$-th user, and $\epsilon$ represents received Gaussian noise with noise power $\sigma^{2}$ and mean 00. The location of the $k$-th user is given by $\bm{p}_{k}=[R_{k},\theta_{k},\phi_{k}]^{T}$, with $R_{k}$, $\theta_{k}$, $\phi_{k}$ being the range, polar angle, and azimuth angle, respectively. Here $\bm{Q}=[\bm{q}_{1},...,\bm{q}_{L_{max}}]\in\mathbb{C}^{3\times L_{max}}$ denotes the locations of the scatters, where the number of scatters is assumed to be less than $L_{max}$, and the location of the $l$-th scatter is given by $\bm{q}_{l}=(r_{l},\vartheta_{l},\varphi_{l})^{T}$. The cascaded channel between the $k$-th user and the BS $\bm{h}_{k}(\bm{p}_{k},\bm{Q})$ is given by to $\bm{h}^{A}_{k}\odot\bm{h}_{k}^{t}(\bm{p}_{k},\bm{Q})$, where $\bm{h}^{A}_{k}$, $\odot$, and $\bm{h}_{k}^{t}(\bm{p}_{k},\bm{Q})$ denote the RIS-BS channel for the $k$-th user, the Hadamard product, and the channel between the RIS and the $k$-th user, respectively. The channel between the $k$-th user and the RIS is given by [14]
$\bm{h}_{k}^{t}(\bm{p}_{k},\bm{Q})=\bm{h}^{t}_{k0}(\bm{p}_{k})+\sum_{l=1}^{L}\bm{h}^{t}_{kl}(\bm{p}_{k},\bm{q}_{l}),$ (2)
where $\bm{h}^{t}_{kl}(\bm{p}_{k},\bm{q}_{l})$ represents the $l$-th path between the RIS and the $k$-th user, and $L$ is the number of scatters in the scenario. We assume $l=0$ is associated with the direct path between the $k$-th user and RIS, while $l\neq 0$ are the paths via scatters. Since users or scatters could be in either NF or FF regions of the RIS, the channels in these two cases are modeled separately.
II-B1 Channel Models for the NF Region
When the user or the scatter is located in the NF region, the signal received by the RIS is described using the spherical wave model and cannot be approximated as a plane wave for the RIS. Thus, the direct user-RIS channel is modeled by[15]
$\bm{h}^{t}_{k0}(\bm{p}_{k})=\alpha_{k}\bm{b}(R_{k},\theta_{k},\phi_{k}),\ \bm{p}_{k}\in\bm{D}_{NF},$ (3)
where $\alpha_{k}$ is the channel gain including the effects of path loss and the directivity of the RIS elements[16], and $\bm{b}$ is the steering vector in the NF region, given as
$\bm{b}(R_{k},\theta_{k},\phi_{k})=\left[\exp\left(-j\dfrac{2\pi}{\lambda}d_{k1}^{t}\right),...,\exp\left(-j\dfrac{2\pi}{\lambda}d_{kN}^{t}\right)\right]^{T}.$ (4)
Here, $d_{kn}^{t}$ is the distance between the $k$-th user and the $n$-th RIS element, and $\bm{D}_{NF}$ represents the NF region. Based on [12], the NF region is defined as the region where the maximum phase error between the phase shift calculated under plane wave approximation and the real phase shift is more than $\pi/8$. Mathematically, the NF region is given by $\bm{D}_{NF}=\{\bm{p}|\Delta\varphi(\bm{p})>{\pi}/{8}\}$, where $\Delta\varphi(\bm{p})$ is the maximum phase error across all the RIS elements, given by
$\Delta\varphi(\bm{p})=\mathop{\max}\limits_{n}\frac{2\pi}{\lambda}\big{(}d_{n}^{t}-(R-y_{n}\sin\theta\sin\phi-z_{n}\cos\theta)\big{)},$ (5)
where $d_{n}^{t}$ is the distance to the $n$-th RIS element, and $(0,y_{n},z_{n})$ is the coordinate of the $n$-th RIS element.
Similarly, we model the paths via the scatters. Specifically, when the $l$-th scatter is located in the NF region, the path between the $l$-th scatter and the $n$-th RIS element is given by
$\bm{h}_{kl}^{t}(\bm{p}_{k},\bm{q}_{l})=\alpha_{kl}\bm{b}(r_{l},\vartheta_{l},\varphi_{l}),\ \bm{q}_{l}\in\bm{D}_{NF},$ (6)
where $\bm{b}(r_{l},\vartheta_{l},\varphi_{l})=[\exp(-j\frac{2\pi}{\lambda}d_{l1}^{t}),...,\exp(-j\frac{2\pi}{\lambda}d_{lN}^{t})]^{T}$, and $d_{ln}^{t}$ is the distance between $l$-th scatter and $n$-th element of the RIS. Here $\alpha_{kl}$ is the channel gain of the $l$-th path.
II-B2 Channel Models for the FF Region
When the user is located in the FF region, we adopt the plane wave approximation to model the received signals. The direct user-RIS channel is modeled by[15]
$\bm{h}^{t}_{k0}(\bm{p}_{k})=\alpha_{k}\bm{a}(\theta_{k},\phi_{k}),\ \bm{p}_{k}\in\bm{D}_{FF},$ (7)
where $\bm{a}(\theta_{k},\phi_{k})$ is the steering vector in the FF region and is given as[17]
$\bm{a}(\theta_{k},\phi_{k})=\left[\exp\left(-j\dfrac{2\pi}{\lambda}(-y_{1}\sin\theta_{k}\sin\phi_{k}-z_{1}\cos\theta_{k})\right),\right.$
$\left....,\exp\left(-j\dfrac{2\pi}{\lambda}(-y_{N}\sin\theta_{k}\sin\phi_{k}-z_{N}\cos\theta_{k})\right)\right]^{T}.$ (8)
In (7), $\bm{D}_{FF}$ represents the FF region, which is defined as the area where the maximum phase error is less than $\pi/8$, i.e., $\bm{D}_{FF}=\{\bm{p}|\Delta\varphi(\bm{p})<\pi/8\}$. When the $l$-th scatter is located in the FF region, the path between the $l$-th scatter and the $n$-th RIS element can be modeled by
$\bm{h}_{kl}^{t}(\bm{p}_{k},\bm{q}_{l})=\alpha_{kl}\bm{a}(\vartheta_{l},\varphi_{l}),\ \bm{q}_{l}\in\bm{D}_{FF}.$ (9)
II-C Localization Protocol
We propose an RIS-enabled hybrid NF and FF source localization protocol, which improves localization accuracy by adaptively localizing the users and optimizing the RIS phase shifts based on the estimated locations. The localization process is divided into $C$ cycles with cycle duration being $\delta$. Each cycle contains three steps: transmission, localization, and optimization. The process of the localization protocol is illustrated in Fig. 2.
II-C1 Transmission
In this step, the users send signals to the RIS for $\delta_{T}$ seconds, which are reflected to the BS. Let $\bm{g}^{(c)}=[g^{(c)}_{1},...,g^{(c)}_{K}]^{T}\in\mathbb{C}^{K\times 1}$ denote the signal received by the BS in the transmission step of the $c$-th cycle. The RIS phase shifts are set randomly in the first cycle, while in the following $C-1$ cycles, the phase shifts are selected based on the optimization results in the previous cycle, which will be described in the optimization step.
II-C2 Localization
In the next $\delta_{E}$ seconds, the BS estimates users’ locations using the received signals. Specifically, in the $c$-th cycle, based on the received signals $\bm{G}^{(c)}=[\bm{g}^{(1)},...,\bm{g}^{(c)}]^{T}\in\mathbb{C}^{K\times c}$, we determine whether the users and the scatters are located in the NF or the FF region and jointly estimate their locations $\bm{P}^{(c)}=[\bm{p}_{1}^{(c)},...,\bm{p}_{K}^{(c)}]$ and $\bm{Q}^{(c)}=[\bm{q}_{1}^{(c)},...,\bm{q}_{L}^{(c)}]$. This step focuses on the problem of accurately locating users given the received signals and the RIS phase shifts. The details of the localization algorithm are introduced in Sec. III.
II-C3 Optimization
In the rest time of the $c$-th cycle, the optimal RIS phase shifts $\bm{\beta}^{(c+1)}$ are selected according to the estimated locations in the former step. Note that this step is not executed in the last cycle. This step focuses on the task of calculating the optimal RIS phase shifts for the next cycle based on the user localization results of the previous step. The details of optimization are introduced in Sec. IV.
In summary, two parts of the high-level problem need to be solved: localization and optimization, which iterate with each other to achieve the optimal RIS phase shifts and higher localization accuracy. Two algorithms are designed to solve these problems respectively.
Figure 2: RIS-enabled hybrid NF and FF source localization protocol.
III RIS-enabled Hybrid Near-field and Far-field Localization
In this section, we first formulate the localization problem and then propose a localization algorithm to solve the formulated problem.
III-A Localization Problem Formulation
We formulate the multi-user localization problem to minimize the user localization loss, which is defined as the sum of the $l_{2}$-norm of the residual between the received signals and the signals reconstructed given user locations. Thus, the localization problem is formulated as
[…]
$s.t.\ \ \bm{p}_{k}\in\bm{D}_{NF}\cup\bm{D}_{FF},\forall k,$ (10b)
where $\bm{g}^{(c)}_{k}=[g_{k}^{(1)},...,g_{k}^{(c)}]^{T}\in\mathbb{C}^{c\times 1}$ is the $k$-th row of received signal $\bm{G}^{(c)}$, i.e. the received signals of the $k$-th user in the previous $c$ cycles. Here $\bm{B}^{(c)}=[\bm{\beta}^{(1)},...,\bm{\beta}^{(c)}]\in\mathbb{C}^{N\times c}$ is the phase shifts of the RIS in the previous $c$ cycles, $\bm{P}=[\bm{p}_{1},...,\bm{p}_{K}]$ is the locations of the users, $\overline{\bm{Q}}$ is the ground truth of scatter locations, and $\|\cdot\|_{2}$ denotes the $l_{2}$-norm. It can be observed from (10a) that the objective function is affected by the accuracy of user location estimation. By solving (P1), we can minimize the estimation error of the user location and eliminate the interference caused by the scattering paths.
Compared to single-user localization, multi-user localization is more challenging because the channels of different users are coupled with each other. Specifically, the signals transmitted by different users at different frequencies may pass through the same scatter. Thus, the localization loss of each user, i.e., a term in the summation (10a), is affected by not only the location of each user but also the scatter locations $\bm{Q}$. This is different from the single-user localization case where the locations of a user and the scatters are optimized to minimize only one localization loss term and the coupling of the localization losses among different users is not considered.
Due to the non-convex nature of the above problem (P1), conventional algorithms like gradient descent can easily fall into local minima. To solve it effectively, a location estimation method by modifying the grid search method is proposed to perform a global search. Specifically, we first sample the search domain in NF and FF regions to create a grid map of the sampled candidate locations, denoted by $\bm{Z}=[\bm{Z}_{near},\bm{Z}_{far}]$, where $\bm{Z}_{near}$ and $\bm{Z}_{far}$ are the sampled candidate locations in the NF and FF, respectively. Then, the most suitable sample locations are selected to minimize the localization loss.
Considering the different signal models in the NF and FF regions, the locations in these two regions are sampled in different ways, which are elaborated in the following.
NF case: In the NF region, $R,\theta,\phi$ are uniformly sampled with sampling spacings $\Delta R,\Delta\theta=\pi/N_{\theta},\Delta\phi=\pi/N_{\phi}$, respectively, where $N_{\theta}$ and $N_{\phi}$ are predetermined parameters. Thus, $S$ sampled candidate locations in the NF are obtained.
$\bm{Z}_{near}=[[R_{1},\theta_{1},\phi_{1}]^{T},...,[R_{S},\theta_{S},\phi_{S}]^{T}].$ (11)
We define the atom channels as the user-RIS channels or the scatter-RIS channels between the RIS and every possible candidate location of the users or the scatters. The atom channel in the NF region is given by[18]
$\bm{F}_{near}=[\bm{b}(R_{1},\theta_{1},\phi_{1}),...,\bm{b}(R_{S},\theta_{S},\phi_{S})]\in\mathbb{C}^{N\times S},$ (12)
where $\bm{b}(R_{i},\theta_{i},\phi_{i})\in\mathbb{C}^{N\times 1}$ is the steering vector given the location $[R_{i},\theta_{i},\phi_{i}]^{T}\in\bm{D}_{NF}$.
FF case: In the FF region, only the angles $\theta,\phi$ are sampled. The range $R$ is not sampled because it does not affect the FF steering vector and it cannot be estimated in a similar method to NF case. The primary reason for this difference lies in the angular relationships between the user and the RIS elements. Specifically, in the NF, the angles between the user and each RIS element are different, leading to different phase shift changes for paths from the user to each RIS element when the range varies. These differences in phase shift changes result in varying received signals, which allows for the estimation of the range $R$. In contrast, for FF users, since the angles between the user and each RIS element are almost the same, when the range changes, the phase shift changes for the path from the user to each RIS element are the same. This uniform phase shift leads to an overall phase change in the received signal, hence range estimation cannot be achieved.
As a result, the number of candidate locations and the algorithm complexity are significantly reduced compared with NF. We apply the same angle sampling methods in the FF as in the NF. The atom channel of the FF region is given by [19]
$\bm{F}_{far}=[\bm{a}(\theta_{1},\phi_{1}),...,\bm{a}(\theta_{N_{\theta}N_{\phi}},\phi_{N_{\theta}N_{\phi}})]\in\mathbb{C}^{N\times N_{\theta}N_{\phi}},$ (13)
where $\bm{a}(\theta_{i},\phi_{i})$ is the steering vector of the FF location $[\theta_{i},\phi_{i}]$.
We define $\bm{F}=[\bm{F}_{near},\bm{F}_{far}]\in\mathbb{C}^{N\times M}$ as the atom channels for the hybrid field, where $M=S+N_{\theta}N_{\phi}$. For better illustration, the sampled locations at $\theta=\pi/2$ plane for both the NF and FF regions are shown in Fig. 3.
Based on the channel model (2) and the atom channels (12) and (13), each user-RIS channel can be approximated as a linear combination of multiple atom channels. We define $\bm{u}_{k}\in\mathbb{C}^{M\times 1}$ as the gains of the atom channels for the $k$-th user-RIS channels, which can be decomposed into two components: $\bm{u}^{\bm{P}}_{k}$ for the direct path, and $\bm{u}^{\bm{Q}}_{k}$ for the scattering path. Both $\bm{u}^{\bm{P}}_{k}$ and $\bm{u}^{\bm{Q}}_{k}$ are unknown, and our aim is to accurately estimate $\bm{u}^{\bm{P}}_{k}$, which contains user locations information. Specifically, $u_{ik}^{\bm{P}}$, the $i$-th element of $\bm{u}^{\bm{P}}_{k}$, denotes the gain of the $i$-th atom channel in the direct path between the $k$-th user and the RIS. If $u_{ik}^{\bm{P}}=0$, it indicates that the $k$-th user is not at location $\bm{z}_{i}$, the ${i}$-th column of the candidate locations $\bm{Z}$. Otherwise, the $k$-th user is located at $\bm{z}_{i}$, and the estimated gain is $u_{ik}^{\bm{P}}$. In contrast, we do not care about the estimation accuracy of $\bm{u}^{\bm{Q}}_{k}$, because it is only influenced by scatter locations.
The $k$-th user-RIS channels are thus given as
$\bm{h}^{t}_{k}(\bm{p}_{k},{\bm{Q}})$ $=\bm{F}\bm{u}_{k}=\bm{F}(\bm{u}^{\bm{P}}_{k}+{\bm{u}^{\bm{Q}}_{k}}).$ (14)
The cascaded channel of the $k$-th user is given by
$\bm{h}_{k}(\bm{p}_{k},{\bm{Q}})=\bm{h}^{A}_{k}\odot\bm{h}_{k}^{t}=\text{diag}\{\bm{h}^{A}_{k}\}\bm{F}(\bm{u}^{\bm{P}}_{k}+{\bm{u}^{\bm{Q}}_{k}}).$ (15)
Note that $\bm{u}_{k}$ can be directly solved base on traditional sparse recovery method like OMP. However, the user location estimated in this way is not accurate enough due to interference of scattering paths. To mitigate this interference and improve user localization accuracy, (P1) is approximated as
[…]
$s.t.$ $\|\bm{u}^{\bm{P}}_{k}\|_{0}=1,\forall k,$ (16b)
where $\|\cdot\|_{0}$ denotes the $l_{0}$-norm, and $\overline{\bm{u}^{\bm{Q}}_{k}}$ is the ground truth of ${\bm{u}^{\bm{Q}}_{k}}$. Here, $\bm{\Lambda}^{k}=(\bm{B}^{(c)})^{T}\text{diag}\{\bm{h}^{A}_{k}\}\bm{F}s_{k}\in\mathbb{C}^{c\times M}$ is the atom signals, which are the linear transformation of the atom channels $\bm{F}$ by considering the effect of RIS phase shift $\bm{B}^{(c)}$, the RIS-BS channel $\bm{h}^{A}_{k}$, and the transmit signal $s_{k}$.
Figure 3: Illustration of the sampled locations at $\theta=\pi/2$.
III-B Localization Algorithm
From (P1’), it is evident that we cannot directly estimate the user’s location since the ground truth of the scatter locations is unknown. Therefore, we need to estimate the locations of the scatters. However, considering that scattering paths are typically weaker than direct user-RIS path, it is difficult to estimate the scatter locations directly from the received signals. To tackle this issue, we leverage the idea of successive interference cancellation (SIC) that first estimates the stronger signal components, i.e., the direct user-RIS paths, and then estimates the weaker scattering paths.
Therefore, we devise a three-step algorithm to solve (P1’). Specifically, we first ignore the scattering path and roughly estimate the locations of the users based on the received signal. The residual signal is then calculated by subtracting the direct user-RIS path component from the received signal to improve the estimation accuracy of scatter locations. Next, the scatter locations are estimated using the residual signal. In the third step, we subtract the scattering path from the received signal and re-estimate the user’s location to achieve higher user localization accuracy. In the following, we introduce these subproblems and the algorithm in detail. The proposed algorithm is summarized in Algorithm 1.
Algorithm 1 The proposed localization algorithm
0:  Recieved signal $\bm{G}^{(c)}$, the atom channels matrix $\bm{F}$, max number of scatters $L_{max}$
0:  The estimated locations of the users $\bm{P}^{(c)}$
1:  Initialization: the atom signals $\bm{\Lambda}^{k}$, the received signals energy $E$, the support set $\mathcal{X}=\emptyset$.
2:  for $k=1,...,K$ do
3:     Obtain the coarse estimated user locations based on (20)
4:     Calculate residual signal by (21) and (22)
5:  end for
6:  Initialize the residual signal $\bm{e}_{k}^{ini}=\bm{e}_{k}$
7:  for $l=1,...,L_{max}$ do
8:     Obtain the estimated $l$-th scatter location by (24)
9:     Update support set $\mathcal{X}=\mathcal{X}\cup i$
10:     Calculate residual signals by (25) and (27)
11:     Calculate the energy of the residual signal $E^{r}$
12:     if $E^{r}<\gamma E$ then
13:        break
14:     end if
15:  end for
16:  for $k=1,...,K$ do
17:     Calculate the new residual signal by (28)
18:     Obtain the estimated location of $k$-th user by (30)
19:  end for
20:  Return: $\bm{P}^{(c)}=[\bm{p}_{1}^{(c)},...,\bm{p}_{K}^{(c)}]$
III-B1 User Location Estimation
In this part, we estimate the location of each user using the corresponding received signal. This is a coarse estimate because we ignore the scattering path in this step. Specifically, we find the atom signal with the highest energy contribution to the received signals and estimate the gain of the atom channel [20] [21]. For example, for the $k$-th user, we use the $\bm{g}_{k}$ to estimate its location, i.e.,
[…]
$s.t.$ $\|\bm{u}^{\bm{P}}_{k}\|_{0}=1,\forall k.$ (17b)
Only one element in $\bm{u}^{\bm{P}}_{k}$ is nonzero, therefore $\bm{u}^{\bm{P}}_{k}$ is a $1$-sparse vector, which enables the sparse recovery method to efficiently recover it[22]. Note that the terms of summation in (17a) are independent, therefore (P1-a) can be decoupled into $K$ independent subproblems, given by
[…]
$s.t.$ $\|\bm{u}^{\bm{P}}_{k}\|_{0}=1.$ (18b)
The received signal $\bm{g}^{(c)}_{k}$ contains multiple paths. Since the gain of the direct path is larger than that of the scattering paths, we choose the atom signal with the strongest correlation with received signal as the direct path corresponding to the user’s location. Mathematically, for the $k$-th user, we have
$i^{k}=\arg\max_{i}|(\bm{\Lambda}^{k}_{i})^{H}\bm{g}^{(c)}_{k}|$ (19)
where $\bm{\Lambda}^{k}_{i}\in\mathbb{C}^{N\times 1}$ is the $i$-th column of $\bm{\Lambda}^{k}$ that corresponds to $\bm{z}_{i}$, the $i$-th location in $\bm{Z}$. Then, we have $\hat{\bm{p}}_{k}=\bm{z}_{i^{k}}$, and the estimated location for all $K$ users is given by
$\hat{\bm{P}}=[\bm{z}_{i^{1}},...,\bm{z}_{i^{K}}].$ (20)
Note that the optimal gain for any given atom signal can be given in closed form as a function $i$
$\hat{u}_{ik}^{\bm{P}}={\bm{\Lambda}^{k}_{i}}^{{\dagger}}\bm{g}^{(c)}_{k},$ (21)
where $\hat{u}_{ik}^{\bm{P}}$ is the estimated value of the $i$-th element in $\bm{u}^{\bm{P}}_{k}$, and  represents the pseudo-inverse. The residual signal for the $k$-th user can be given by
$\bm{e}^{(c)}_{k}=\bm{g}^{(c)}_{k}-\hat{u}_{ik}^{\bm{P}}\bm{\Lambda}^{k}_{i}.$ (22)
In this way, signals with smaller interference for scatter localization is obtained.
III-B2 Scatter Location Estimation
In this part, the locations of the scatters are estimated using all the residual signals. This part is iteratively conducted until the energy of the new residual signal is less than a percentage of the energy of the received signals or the number of iterations reaches $L_{max}$. This subproblem is modeled as
[…]
$s.t.$ |$|\{i|\bm{v}_{i}\neq\bm{0}\}|\leq L_{max},$ (23b)
where $\bm{v}_{i}=[u_{i1}^{\bm{Q}},...,u_{iK}^{\bm{Q}}]\in\mathbb{C}^{1\times K}$. Since the numbers of scatters are generally much smaller than the number of atom channels[23], we can use the sparse recovery method[20] to estimate the sparse vector $\bm{u}^{\bm{Q}}_{k}$[22]. We define the support set $\mathcal{X}$ as the indices of the atom signals corresponding to the estimated scatter locations. In each iteration, a new index is added to the support set. Note that the atom signal for each user is different, the residual signal for different users is approximated using their corresponding atom signals, and then the location with the minimal modulus sum of the new residual signals is selected as the estimated location of a scatter. Mathematically, we have
$i=\arg\max_{i}\sum_{k=1}^{K}|(\bm{\Lambda}^{k}_{i})^{H}\bm{e}^{(c)}_{k}|$ (24)
Then, we update the support set as $\mathcal{X}=\mathcal{X}\cup i$. The gains of all the support signals in the support set are calculated through orthogonal least square, which is given by
$\hat{\bm{u}}^{\bm{Q}}_{k,\mathcal{X}}=({\bm{\Lambda}^{k}_{\mathcal{X}}})^{{\dagger}}\bm{e}_{k}^{ini},$ (25)
where $\bm{e}_{k}^{ini}$ is the initial residual signal, and $\bm{\Lambda}_{\mathcal{X}}^{k}$ is the matrix generated by the atom signals selected by the indices in the support set, which is given by
$\bm{\Lambda}^{k}_{\mathcal{X}}=[\bm{\Lambda}^{k}_{j}],j\in\mathcal{X}.$ (26)
Using the estimated gain, the residual signal is updated by
$\bm{e}^{(c)}_{k}=\bm{e}_{k}^{ini}-\bm{\Lambda}^{k}_{\mathcal{X}}\hat{\bm{u}}^{\bm{Q}}_{k,\mathcal{X}}.$ (27)
Then the energy of the new residual signal $\bm{e}_{k}$ is calculated and denoted by $E^{r}$. If $E^{r}<\gamma E$, the iteration terminates, else the steps (24) to (27) are repeated, where $\gamma$ is a predetermined parameter and $E$ is the energy of the original received signals.
III-B3 User Location Refinement
In this part, we optimize the locations of users based on the estimated scatter locations, i.e., re-solve the locations of users given the locations of scatters. First, we calculate the residual signal $\bm{r}^{(c)}_{k}$ by subtracting the scattering path from the original received signal.
$\bm{r}^{(c)}_{k}=\bm{g}^{(c)}_{k}-\bm{\Lambda}^{k}_{\mathcal{X}}{\bm{\Lambda}^{k}_{\mathcal{X}}}^{{\dagger}}\bm{g}^{(c)}_{k}.$ (28)
Then similar to the method of the user location estimation step, by solving the following subproblem,
[…]
$s.t.$ $\|\bm{u}^{\bm{P}}_{k}\|_{0}=1,\forall k,$ (29b)
the index of the refined $k$-th user location is given by
$i^{k}=\arg\max_{i}|(\bm{\Lambda}^{k}_{i})^{H}\bm{r}^{(c)}_{k}|.$ (30)
Then we have $\hat{\bm{p}}_{k}=\bm{z}_{i^{k}}$ and $\hat{\bm{P}}=[\bm{z}_{i^{1}},...,\bm{z}_{i^{K}}]$.
IV RIS Phase Shift Optimization
In this section, we first formulate the RIS phase shift optimization problem and propose an optimization algorithm to solve it.
IV-A Phase Shift Optimization Problem Formulation
The selected optimization metric is CRB, a standard metric for assessing the parameter estimation error. The CRB is derived from the model of the received signal for $c+1$ cycles and given by
$\bm{y}_{k}=[y^{(1)}_{k},...,y^{(c+1)}_{k}]^{T}=(\bm{B}^{(c+1)})^{T}\bm{h}_{k}s+\bm{\epsilon},$ (31)
where $\bm{\epsilon}=[\epsilon^{(1)},...,\epsilon^{(c+1)}]^{T}$ is the independent and identically distributed zero-mean Gaussian noise with variance $\sigma^{2}$. Note that the received signal is influenced by the RIS phase shift $\bm{\beta}^{(c+1)}$, which is the $(c+1)$-th column of $\bm{B}^{(c+1)}$. Thus, we can obtain the expressions between $\bm{\beta}^{(c+1)}$ and CRB, listed in the following. By optimizing the RIS phase shifts $\bm{\beta}^{(c+1)}$, we can minimize the CRB to improve the localization accuracy.
Specifically, we formulate the optimization problem to minimize the sum of weighted CRBs of the range and angle estimation errors. Note that we estimate the angles and ranges for NF users, while only the angles for FF users. In other words, the location parameters for NF and FF users are different. To this end, we give the expressions of the weighted CRBs for NF and FF users separately.
IV-A1 CRBs for NF Users
If the estimated location $\bm{p}_{k}^{(c)}$ is in the NF region in the $c$-th cycle, we expect that in the $(c+1)$-th cycle, the range and angles of the $k$-th user are estimated, and the estimated parameters are denoted by $\bm{p}_{k}^{(c+1)}=[R_{k},\theta_{k},\phi_{k}]^{T}$. The expressions of CRBs for the NF case are provided in proposition IV-A1.
Proposition 1: For the NF case, the CRBs of the unknown parameters $\bm{p}_{k}=[R_{k},\theta_{k},\phi_{k}]^{T}$ are given by
$D_{R}$ $=\left(\bm{J}_{NF}^{-1}\right)_{1,1},$ (32)
$D_{\theta}$ $=\left(\bm{J}_{NF}^{-1}\right)_{2,2},$ (33)
$D_{\phi}$ $=\left(\bm{J}_{NF}^{-1}\right)_{3,3},$ (34)
where $\bm{J}_{NF}$ is the $3\times 3$ Fisher information matrix of $\bm{p}_{k}$ in the NF region, and the $(i,j)$-th element in $\bm{J}_{NF}$ is given by [24]
$[\bm{J}_{NF}]_{i,j}=\dfrac{2}{\sigma^{2}}\sum_{m=1}^{c+1}Re\left\{\dfrac{\partial(\mu^{(m)})^{H}}{\partial p_{i}}\dfrac{\partial\mu^{(m)}}{\partial p_{j}}\right\},$ (35)
where $\mu^{(m)}$ is defined as the noise-free received signal
$\mu^{(m)}=(\bm{\beta}^{(m)})^{T}\bm{h}^{NF}_{k}s_{k}.$ (36)
Here $\bm{h}^{NF}_{k}$ is the channel calculated by using $\bm{p}_{k}^{(c)}\in\bm{D}_{NF}$ and $\bm{Q}^{(c)}$, and $p_{i}$ is the $i$-th element in $\bm{p}_{k}$.
Proof:
See Appendix A. ∎
IV-A2 CRBs for FF Users
If the estimated location $\bm{p}_{k}^{(c)}$ is in the FF region, we expect that only the angles are estimated for the $k$-th user in the $(c+1)$-th cycle, and the estimated parameters are denoted by $\bm{p}_{k}^{(c+1)}=[\theta_{k},\phi_{k}]^{T}$. The expressions of CRBs for the FF case are provided in proposition IV-A2.
Proposition 2: For the FF case, the CRBs of the unknown parameters $\bm{p}_{k}$ are given by
$D_{\theta}$ $=\left(\bm{J}_{FF}^{-1}\right)_{1,1},$ (37)
$D_{\phi}$ $=\left(\bm{J}_{FF}^{-1}\right)_{2,2},$ (38)
where $\bm{J}_{FF}$ is the $2\times 2$ Fisher information matrix of $\bm{p}_{k}$ in the FF region, and the $(i,j)$-th element is given by [24]
$[\bm{J}_{FF}]_{i,j}=\dfrac{2}{\sigma^{2}}\sum_{m=1}^{c+1}Re\left\{\dfrac{\partial(\mu^{(m)})^{H}}{\partial p_{i}}\dfrac{\partial\mu^{(m)}}{\partial p_{j}}\right\},$ (39)
where $\mu^{(m)}$ is defined as the noise-free received signal
$\mu^{(m)}=(\bm{\beta}^{(m)})^{T}\bm{h}^{FF}_{k}s_{k},$ (40)
and $\bm{h}^{FF}_{k}$ is the channel calculated using $\bm{p}_{k}^{(c)}$ and $\bm{Q}^{(c)}$.
Proof:
See Appendix B. ∎
IV-A3 Optimization Problem Formulation
To achieve high localization accuracy for multiple users, we optimize the phase shifts $\bm{\beta}^{(c+1)}$ by minimizing the sum of CRBs of all the users. The optimization problem can be formulated as
[…]
$+\sum_{k\in\Psi_{FF}}tr\left(\bm{J}_{FF,k}^{-1}\bm{W}_{FF}\right),$ (41a)
$s.t.\ |$ $\beta_{n}^{(c+1)}|=1,\ \forall n=1,2,...N,$ (41b)
where $\bm{J}_{NF,k}$ or $\bm{J}_{FF,k}$ is the Fisher information matrix of $\bm{p}_{k}$ when the $k$-th user is located in the NF or FF region and is a function of the RIS phase shifts $\bm{\beta}^{(c+1)}$. Here $\bm{W}_{NF}=\text{diag}\{w_{1},w_{2},w_{3}\}$ and $\bm{W}_{FF}=\text{diag}\{w_{4},w_{5}\}$ are the weight matrices of the CRBs in the NF and FF regions, respectively. $\Psi_{NF}$ and $\Psi_{FF}$ are the sets of estimated NF and FF users, which are mathematically given by
$\Psi_{NF}=\{k|\bm{p}_{k}^{(c)}\in\bm{D}_{NF}\},$ (42)
$\Psi_{FF}=\{k|\bm{p}_{k}^{(c)}\in\bm{D}_{FF}\}.$ (43)
$\beta_{n}^{(c+1)}$ is the RIS phase shift of the $n$-th element in the $(c+1)$-th cycle, satisfying the constant modulus constraint (41b).
IV-B RIS Phase Shift Optimization Algorithm
We design an optimization algorithm based on the CCM method to tackle (P2). Due to the constant modulus constraint (41b), the optimization problem (P2) is non-convex, which is numerically difficult to handle. Fortunately, the solution can be considered as lying on the CCM to satisfy the constant modulus, where the manifold is represented as
$\mathcal{M}^{N}=\left\{\bm{\beta}^{(c+1)}\in\mathbb{C}^{N}:|\beta_{1}^{(c+1)}|=...=|\beta_{N}^{(c+1)}|=1\right\}.$ (44)
The main idea of the CCM-based optimization method is to iteratively apply gradient descent in the manifold space. After several iterations, the algorithm terminates when the difference between two iterations of $f(\bm{P},\bm{\beta}^{(c+1)})$ is less than a constant $\zeta$. The CCM-based optimization method is composed of four main steps in each iteration:
IV-B1 Compute the gradient in Euclidean space
We use the Euclidean gradient as the search direction for the minimization problem in Euclidean space. The Euclidean gradient of $f(\bm{\beta}^{(c+1)})$ is given by
$\nabla f(\bm{\beta}^{(c+1)})=2\Big{(}$ $\sum_{k\in\Psi_{NF}}\left[\dfrac{w_{1}\partial D_{R,k}}{\partial{\bm{\beta}^{(c+1)}}^{\ast}}+\dfrac{w_{2}\partial D_{\theta,k}}{\partial{\bm{\beta}^{(c+1)}}^{\ast}}+\dfrac{w_{3}\partial D_{\phi,k}}{\partial{\bm{\beta}^{(c+1)}}^{\ast}}\right]$
$+$ $\sum_{k\in\Psi_{FF}}\left[\dfrac{w_{4}\partial D_{\theta,k}}{\partial{\bm{\beta}^{(c+1)}}^{\ast}}+\dfrac{w_{5}\partial D_{\phi,k}}{\partial{\bm{\beta}^{(c+1)}}^{\ast}}\right]\Big{)},$ (45)
where the variable $\bm{P}$ is omitted in $f(\bm{P},\bm{\beta}^{(c+1)})$ for simplicity. Similar to the proof in [25], the specific expressions of the differentials of CRBs are given in Appendix C.
IV-B2 Compute the Riemannian gradient
The Riemannian gradient is the projection of the Euclidean gradient onto the tangent space of the complex circle manifold. The Riemannian gradient of the objective function $f(\bm{\beta}^{(c+1)})$ at the point $\bm{\beta}^{(c+1)}_{j}$ on the complex circle manifold $\mathcal{M}$ is given as [25]
$\nabla_{\mathcal{M}}$ $f(\bm{\beta}^{(c+1)}_{j})=-\nabla f(\bm{\beta}^{(c+1)}_{j})$
$-Re\left\{\left(\nabla f(\bm{\beta}^{(c+1)}_{j})\right)^{*}\odot{\bm{\beta}^{(c+1)}_{j}}\right\}\odot\bm{\beta}^{(c+1)}_{j}.$ (46)
IV-B3 Update over the Tangent Space
We choose a step size to update the current point, which is mathematically given as
$\bar{\bm{\beta}}^{(c+1)}_{j}=\bm{\beta}^{(c+1)}_{j}+\alpha_{j}\nabla_{\mathcal{M}}g(\bm{\beta}^{(c+1)}_{j}),$ (47)
where $\alpha_{j}$ is the step size in the $j$-th iteration.
IV-B4 Retract onto the manifold
After the update, the new point $\bar{\bm{\beta}}^{(c+1)}_{j}$ generally does not lie on the manifold $\mathcal{M}$. By using the retraction operator, the new point is mapped into the manifold. The retraction operator is given as
$\bm{\beta}^{(c+1)}_{j+1}=\bar{\bm{\beta}}^{(c+1)}_{j}\odot\dfrac{1}{|\bar{\bm{\beta}}^{(c+1)}_{j}|}.$ (48)
The proposed algorithm is summarized in Algorithm 2.
Algorithm 2 CCM-based RIS Phase Shift Optimization Algorithm
0:  Estimated location $\hat{\bm{p}}^{(c)}$, RIS phase shift for previous $c$ cycles $\bm{B}^{(c)}$
0:  Optimal RIS phase shift $\bm{\beta}^{(c+1)}$
1:  Initialize: $j=0$, $\beta_{0}\in\mathcal{M}$;
2:  while $|g(\bm{\beta}^{(c+1)}_{j+1})-g(\bm{\beta}^{(c+1)}_{j})|<\zeta$ do
3:     Compute the Euclidean gradient $\nabla f(\bm{\beta}^{(c+1)}_{j})$ according to (IV-B1)
4:     Calculate the Riemannian gradient $\nabla_{\mathcal{M}}f(\bm{\beta}^{(c+1)}_{j})$ according to (IV-B2);
5:     Compute the RIS phase shift update on the tangent space $\bar{\bm{\beta}}^{(k+1)}_{j}$ according to (47);
6:     Update RIS phase shift $\bm{\beta}^{(k+1)}_{j+1}$ according to (48);
7:     $j=j+1$;
8:  end while
V Performance Analysis
In this section, we analyze the complexity of the proposed method and discuss its localization performance.
V-A Algorithm Complexity
V-A1 Complexity of the localization algorithm
For the user location estimation step of the localization algorithm, the method is similar with the first iteration of the OMP method. Therefore, the computational cost of the user estimation step is $\mathcal{O}(cMK)$[26]. In the scatter location estimation step, the complexity of this algorithm has been analyzed in [27]. We estimate the scatter locations for at most $L_{max}$ rounds. Hence, the computational cost is $\mathcal{O}(cKML_{max})$. Then, similar with the user location estimation step, the computational cost of user location refinement step is $\mathcal{O}(cMK)$.
Hence, the overall cost of the localization algorithm for the $c$-th cycle is $\mathcal{O}(cMK)+\mathcal{O}(cKML_{max})$. The overall complexity for a $C$-cycle localization algorithm is $\mathcal{O}(C(C+1)MK)+\mathcal{O}(C(C+1)KML_{max})$.
V-A2 Complexity of the RIS optimization algorithm
The complexity of the CCM algorithm has been analyzed in [28]. Let us denote the total number of iterations required to converge by $T_{CCM}$. Then the total complexity of the optimization algorithm is $\mathcal{O}(T_{CCM}N^{2})$. Since the optimization algorithm is conducted for $C-1$ times, the overall complexity for the optimization algorithm is $\mathcal{O}((C-1)T_{CCM}N^{2})$.
V-B Localization Performance
Localization errors primarily arise from two sources: one is the failure to estimate the grid closest to the user, which we refer to as grid misjudgment; the other is the error between the user’s continuous location in actual space and the discrete grids, known as grid mismatch, reflecting the inherent limitation of discretizing continuous space[29].
Mathematically, suppose the real location of the $k$-th user is $\bm{p}_{k}$, the estimation result is $\bm{z}_{m^{\prime}}$, and the nearest grid for the $k$-th user is $\bm{z}_{m}$. Then the average localization error can be given by
$\mathbb{E}(l_{e})$ $=\mathbb{E}\{\|\bm{p}_{k}-\bm{z}_{m^{\prime}}\|\}$
$\leq\mathbb{E}\{\|\bm{p}_{k}-\bm{z}_{m}\|\}+\mathbb{E}\{\|\bm{z}_{m}-\bm{z}_{m^{\prime}}\|\},$ (49)
where $l_{e}=\|\bm{p}_{k}-\bm{z}_{m^{\prime}}\|$ is the localization error. The first term corresponds to grid mismatch, and the second term is grid misjudgment. In the following, we analyze these two sources of errors respectively.
V-B1 Number of Cycles
The expectation of localization error can be given by[30]
[…]
where $\mathcal{R}_{k,m}$ is the decision region for the $m$-th candidate location. The integral in (50) is the probability that the estimated location is the $m$-th candidate location, given the ground truth that the location of the $k$-th user is $\bm{p}_{k}$. Here $\gamma_{m}^{k}$ is the error parameter, defined as $\gamma_{m}^{k}=\|\bm{p}_{k}-\bm{z}_{m}\|$.
We assume all locations have the same prior probabilities. Hence, the decision region can be given by
$\mathcal{R}_{k,m}=\{\bm{g}_{k}^{(C)}:$ $\mathbb{P}(\bm{g}_{k}^{(C)}|\bm{B}^{(C)},\bm{z}_{m},\bm{Q})$
$\geq\mathbb{P}(\bm{g}_{k}^{(C)}|\bm{B}^{(C)},\bm{z}_{m^{\prime}},\bm{Q}),\forall m^{\prime}\neq m\}.$ (51)
Since the noise obeys Gaussian distribution, we have
$\mathcal{R}_{k,m}=\{\bm{g}_{k}^{(C)}:|\bm{g}_{k}^{(C)}-\bm{\mu}_{m}|^{2}\leq|\bm{g}_{k}^{(C)}-\bm{\mu}_{m^{\prime}}|^{2},\forall m^{\prime}\neq m\},$ (52)
where $\bm{\mu}_{m}=[\mu^{(1)}_{m},...,\mu^{(C)}_{m}]^{T}$ is the noise-free signal for the $m$-th grid. Let $\bm{\xi}_{k}=\bm{g}_{k}^{(C)}-\bm{\mu}_{m}$, we have[30]
$\mathcal{R}_{k,m}=$$\{\bm{g}_{k}^{(C)}:|\bm{\xi}_{k}|^{2}\leq|\bm{\xi}_{k}+\bm{\mu}_{m}-\bm{\mu}_{m^{\prime}}|^{2},\forall m^{\prime}\neq m\},$
$=$ $\{\bm{g}_{k}^{(C)}:|\bm{\mu}_{m}-\bm{\mu}_{m^{\prime}}|^{2}+2\bm{\xi}_{k}(\overline{\bm{\mu}_{m}-\bm{\mu}_{m^{\prime}}})\geq 0,$ (53)
$\ \forall m^{\prime}\neq m\}.$
Assuming the user is located at a grid location $\bm{z}_{m}$, when $C\to\infty$, the first term is greater than 0, and the second term converges to 00 because $\bm{\xi}_{k}$ and $\bm{\mu}_{m}-\bm{\mu}_{m^{\prime}}$ are independent and $\mathbb{E}(\bm{\xi}_{k})=0$. Therefore, we have
$\lim_{C\to\infty}\mathbb{P}(\bm{g}_{k}^{(C)}\in\mathcal{R}_{k,m})=1.$ (54)
Hence, when ignoring grid mismatch, the grid misjudgment error $\mathbb{E}\{\|\bm{z}_{m}-\bm{z}_{m^{\prime}}\|\}$ converges to 00.
When the user is not at a grid, if we assume the midpoint of $\bm{z}_{m}$ and $\bm{z}_{m^{\prime}}$ also corresponds the midpoint of $\bm{\mu}_{m}$ and $\bm{\mu}_{m^{\prime}}$, then we have $\mathbb{E}(\bm{\xi}_{k})\leq|\bm{\mu}_{m}-\bm{\mu}_{m^{\prime}}|/2$. Hence, the $|\bm{\mu}_{m}-\bm{\mu}_{m^{\prime}}|^{2}+2\bm{\xi}_{k}(\overline{\bm{\mu}_{m}-\bm{\mu}_{m^{\prime}}})\geq 0$ is still alwasy true, and we still have $\lim_{C\to\infty}\mathbb{P}(\bm{g}_{k}^{(C)}\in\mathcal{R}_{k,m})=1$. However, the number of cycles does not affect the grid mismatch error. Therefore, the localization error would converge to a non-zero number.
Remark 1: By increasing the number of estimation cycles $C$, the localization error first decreases and then remains fixed.
V-B2 Sampling Spacing
First, we consider the scenario $C\to\infty$. According to the previous analysis, the localization error
$\mathbb{E}(l_{e})$ $\leq\mathbb{E}\{\|\bm{p}_{k}-\bm{z}_{m}\|\}+\mathbb{E}\{\|\bm{z}_{m}-\bm{z}_{m^{\prime}}\|\},$
$\overset{C\to\infty}{\approx}\mathbb{E}\{\|\bm{p}_{k}-\bm{z}_{m}\|\},$ (55)
where $\mathbb{E}\{\|\bm{p}_{k}-\bm{z}_{m}\|\}$ is directly related to the sampling spacing. Hence, a finer grid sampling could alleviate the grid mismatch problem[29].
When constrained by a limited number of cycle, the approximation in (V-B2) no longer holds. Note that $\bm{\xi}_{k}=\bm{g}_{k}^{(C)}-\bm{\mu}_{m}$ in (V-B1) follows Gaussian distribution. At a given SNR, decreasing the sampling spacing raises the probability that $|\bm{\xi}_{k}|>|\bm{\mu}_{m}-\bm{\mu}_{m^{\prime}}|/2$. This leads to a reduction in the probability of correctly estimating the nearest grid and an increase in the expected error between the estimated and nearest grid points, i.e. $\mathbb{E}\{\|\bm{z}_{m}-\bm{z}_{m^{\prime}}\|\}$. Furthermore, the effectiveness of sparse recovery methods is constrained by the restricted isometry property (RIP) or the mutual coherence condition. Reducing the sampling spacing may breach this condition, leading to a decline in localization accuracy[31].
V-B3 Number of RIS Elements
Similarly to [32], we characterize the average received power with respect to the number of RIS elements as $N\to\infty$. We consider two different RIS phase shifts: random phase shifts and optimal phase shifts that maximize the SNR. We consider $C=1$ as an example.
Proposition 3: Assume $\bm{h}_{k}^{A}\sim\mathcal{CN}(\bm{0},\rho_{A}^{2}\bm{I})$ and $\bm{h}_{k}^{t}\sim\mathcal{CN}(\bm{0},\rho_{t}^{2}\bm{I})$, the average received power holds that
$\lim_{N\to\infty}P=\begin{cases}N\rho_{A}^{2}\rho_{t}^{2},&\bm{\beta}=[1,...,1]^{T},\\ N^{2}\dfrac{\pi^{2}\rho_{A}^{2}\rho_{t}^{2}}{16},&\text{max SNR phase shifts}.\end{cases}$ (56)
Proof:
The two cases are discussed as follows:
  • The channel $h=(\bm{h}^{A}_{k})^{H}\text{diag}\{\bm{\beta}\}\bm{h}_{k}^{t}=(\bm{h}^{A}_{k})^{H}\bm{h}_{k}^{t}$. According to Lindeberg-Lévy central limit theorem, we have $h\sim\mathcal{CN}(\bm{0},N\rho_{A}^{2}\rho_{t}^{2})$ as $N\to\infty$. The average user received power is given by
    $\lim_{N\to\infty}P=\mathbb{E}|h|^{2}=N\rho_{A}^{2}\rho_{t}^{2}.$ (57)
  • For RIS phase shifts of max SNR, we have $|h|=(\bm{h}^{A}_{k})^{H}\text{diag}\{\bm{\beta}\}\bm{h}_{k}^{t}=\sum_{n=1}^{N}|h^{A}_{k,n}||h_{k,n}^{t}|$, where $h^{A}_{k,n}$ and $h_{k,n}^{t}$ are the $n$-th element in $\bm{h}^{A}_{k}$ and $\bm{h}_{k}^{t}$, respectively. Since $h^{A}_{k,n}$ and $h_{k,n}^{t}$ are statistically independent and follow Rayleigh distribution with mean values $\sqrt{\pi}\rho_{A}/2$, $\sqrt{\pi}\rho_{t}/2$, we have $\mathbb{E}(|h^{A}_{k,n}||h_{k,n}^{t}|)=\pi\rho_{A}\rho_{t}/4$. The average user received power is given by
    $\lim_{N\to\infty}P$ $=\lim_{N\to\infty}\left|\sum_{n=1}^{N}|h^{A}_{k,n}||h_{k,n}^{t}|\right|^{2}=N^{2}\dfrac{\pi^{2}\rho_{A}^{2}\rho_{t}^{2}}{16}.$ (58)
Note that the power gain of order $N$ can be achieved for fixed phase shifts, which reveals the inherent aperture gain of a larger RIS by collecting more signal power. Moreover, setting the RIS phase shifts for max SNR can also achieve another beamforming power gain of order $N$ simultaneously.
Hence, with an increasing number of RIS elements, the SNR improves, which enlarges the difference between $\bm{\mu}_{m}$ and $\bm{\mu}_{m^{\prime}}$. Consequently, the probability of $|\bm{\xi}_{k}|>|\bm{\mu}_{m}-\bm{\mu}_{m^{\prime}}|/2$ decreases, which lowers the likelihood of grid misjudgment. However, the impact of grid mismatch persists.
Remark 2: As the number of RIS elements increases, the localization accuracy first improves and then remains fixed.
VI Simulation Results
In this section, we present simulation results demonstrating the performance of the proposed method. Simulations are conducted on a personal computer with a 1.80GHz Intel Core i7 processor. The RIS is placed on the plane $x=0$, and the center of RIS is at the origin $(0,0,0)$m. We use an RIS with $10\times 10$ elements, and the element spacing is $\lambda/2=0.005$m. The center transmit frequency is 28GHz, and the frequency spacing between different users is set as $\Delta f=120$kHz[33]. The localization range is $10$m in radius. The azimuth and elevation angles are both uniformly sampled with sampling spacing $\pi/10$, and the range is sampled with spacing $0.25$m. The number of cycles is set as $20$. We conduct $T=1000$ independent trials to obtain the average result. The simulation parameters are listed in Table I unless otherwise stated.
TABLE I: Simulation Parameters
Parameters Values
Number of users $K$ $2$
Number of scatters $L$ $3$
Transmit center frequency $f_{c}$ $28$GHz
Frequency spacing $\Delta f$ $120$KHz
RIS center location $(0,0,0)$m
Number of RIS element $N$ $10\times 10$
RIS element spacing 0.005m
Localization range 10m
Angle sampled spacing $\Delta\theta,\Delta\phi$ $\pi/10$ rad
Range sampled spacing $\Delta R$ i 0.25m
Number of cycles $C$ 20
VI-A Performance Comparison
Figure 4: (a) Angle estimation accuracy of the proposed algorithm compared with AML and ANM when the user is located in the FF region. (b) Angle estimation accuracy of the proposed algorithm compared with AML and ANM when the user is located in the NF region. (c) Range estimation accuracy of the proposed algorithm compared with AML when the user is located in the NF region.
Fig. 4 shows the RMSE of angle or range estimation versus the signal-to-noise ratio (SNR) when the user is in the FF or the NF region. The SNR is defined as $P_{s}/\sigma^{2}$, where $P_{s}=((\bm{h}^{A}_{k})^{T}\bm{I}\bm{h}_{k0}^{t})^{2}$ is the received signal power of the direct path when the phase shifts of all the RIS elements are $1$. The proposed method is compared with two kinds of comparison methods.
  • First, the results of a RIS-enabled NF localization method (AML in [16]) and a RIS-enabled FF localization method (ANM in [34]) are provided for comparison to reveal the advantage of the proposed algorithm.
  • Second, the proposed algorithm is combined with pure NF or FF models to show the gain brought by the hybrid model, labeled as Proposed (NF) and Proposed (FF).
Fig. 4 (a) shows the angle estimation RMSE of the proposed algorithm and other comparison algorithms when the user is located in the FF region. Compared with AML and ANM, it shows that the proposed method outperforms all other comparison methods for FF user localization. We also observe that for algorithms other than ANM, as SNR increases, the RMSE of the angle estimation first decreases and then remains fixed. The RMSE does not decrease when the SNR is very large because the estimated location must be at one of the sampled grids, while the actual location of the users could be off-grid. The ANM does not suffer from the grid mismatch problem since it is an off-grid algorithm.
Compared with the results of pure NF or FF models, Fig. 4 (a) reveals that the RMSEs of the proposed algorithm when applied to the hybrid model and FF model are similar, indicating the hybrid model behaves similarly for FF user localization although with the interference of NF candidate locations. We also observe that the proposed algorithm performs worse in the NF model than in the hybrid model for FF user localization. This is because, for the pure NF model, the range of the FF region is also sampled, resulting in more candidate locations in the FF region and making the algorithm more difficult to converge. At the same time, the channels of the sampled FF candidate locations in the NF model have strong correlations, which leads to energy dispersion especially when estimating the scatters locations, resulting in lower estimation performance. We conducted more simulations in the following to support the above arguments.
Fig. 4 (b) shows the angle estimation RMSE of the proposed algorithm when the user is located in the NF region. It shows that the proposed method provides higher localization accuracy than the compared algorithms. We also observe that the FF model has a higher RMSE than the hybrid model when estimating the NF angles. This is due to the model mismatch of the FF model when describing the NF channels. Therefore, the hybrid model has the advantage of higher localization accuracy compared to the pure FF model. Furthermore, for NF user localization, the angle estimation result of the proposed algorithm applied to the hybrid and NF models is similar.
Fig. 4 (c) shows the range estimation accuracy for NF users versus SNR. For range estimation, the proposed method outperforms AML by more than 55% for all SNR conditions. It shows that the range estimation result of the NF model is not as good as those of the hybrid model for NF user localization, which reveals the superiority of the hybrid model.
Figure 5: Angle estimation performance versus number of cycles for FF users.
Fig. 5 illustrates the angle estimation RMSE versus the number of cycles for FF users. It shows that for the hybrid model and FF model, the estimation performance converges after $C=30$ cycles, while for the NF model, the result remains fixed after $C=35$ cycles, which proves the argument that the NF model is harder to converge.
Figure 6: Angle estimation performance versus localization ranges for FF users.
Fig. 6 shows the angle estimation RMSE versus the localization range. We observe that as the localization range expands, the RMSE under the NF model increases, while the RMSE remains unchanged for the hybrid model. This is because as the localization range expands, for the NF model, the number of candidate locations also increases, while they remain the same for the FF and hybrid model. Also, note that for a sufficiently small estimation range, the NF model could outperform the hybrid model. This is due to that the NF model is more accurate and the FF model is the approximation of the NF model for the FF region. When the localization range is sufficiently small, the above issue of the NF model is mitigated, and the advantage of the NF model being more accurate exceeds the losses caused by the above problem.
Figure 7: CPU time of the proposed algorithm versus number of cycles.
Fig. 7 reveals the CPU time versus number of cycles for three different models. The results indicate that the proposed algorithm has the lowest complexity when applied to the FF model, followed by the hybrid model, and the highest complexity in the NF model. This is due to that the FF model has the fewest candidate locations, while the NF model has the most. This shows the advantage of low algorithm complexity of the hybrid model over the NF model.
In conclusion, the proposed method outperforms other RIS-based localization algorithms in both angle and range estimation accuracy. Compared with different models, the hybrid model offers superior localization accuracy compared to the pure FF model and has the dual advantage of higher localization accuracy and lower algorithm complexity compared to the pure NF model.
VI-B Localization Accuracy and Complexity Trade-off
Figure 8: Average angle estimation performance of the NF and FF users for different angle sampling spacing.
Figure 9: CPU time of the proposed algorithm versus angle sampling spacing.
This subsection examines the effects of hyperparameters on localization accuracy and algorithm complexity. The analysis in Sec. V shows that the sampling spacing, the number of cycles, and the number of RIS elements are the most important factors for both.
VI-B1 Number of Cycles
Fig. 5 and Fig. 7 illustrate the angle estimation RMSE and the CPU time for different numbers of cycles. The RMSE is negatively correlated with the number of cycles, while the CPU time is positively correlated with the number of cycles, which demonstrates the trade-off relationship between localization accuracy and algorithm complexity.
VI-B2 Sampling Spacing
Fig. 8 shows the average angle estimation RMSE of the proposed algorithm under different sampling spacing. Fig. 9 shows the CPU time of the proposed algorithm in three different models. It can be observed that as the angle sampling spacing reduces, the CPU time increases rapidly, while the estimation error first decreases and then remains fixed, which is consistent with previous analysis. Note that when the spacing is less than $\pi/40$, the accuracy does not improve, but the complexity still increases.
VI-B3 RIS Size
We set the number of cycles as $C=50$ to ensure convergence. Fig. 1 and Fig. 11 show the range estimation RMSE and the CPU time of the proposed algorithm versus the number of RIS elements. As the number of RIS elements increases, the range estimation RMSE decreases, and CPU time increases. This is because the complexities of both the localization and optimization algorithms are positively correlated with the RIS element number. Also, a centimeter-level accuracy is attainable for a RIS with $900$ elements.
Figure 10: Range estimation performance of the NF users versus RIS size.11footnotemark: 1
Figure 11: CPU time of the proposed algorithm for different RIS sizes.
VII Conclusion
In this paper, a RIS-enabled multi-user hybrid NF and FF localization method has been developed. We considered a multi-user hybrid NF and FF localization scenario and introduced a localization protocol. A localization algorithm is designed for location estimation. The CRB of the estimated location parameters was analyzed and a CCM-based optimization method is proposed to optimize RIS phase shifts. We analyzed the complexity and the localization performance of the proposed method. Simulation results have shown that: 1) the proposed method outperforms the compared algorithm and can achieve sub-meter accuracy; 2) the hybrid NF and FF model has the advantages of high accuracy and low algorithm complexity over pure NF and FF models; 3) by adjusting the hyperparameters, the trade-off between localization accuracy and complexity can be achieved.
Appendix A Proof of Proposition IV-A1
As mentioned in Section III-A, all previously received signals are used to estimate the locations, so we need to consider received signals in all previous $c$ cycles, as modeled in (31), to calculate the CRBs.
To obtain the expressions of CRBs, we first derive the expression of the Fisher information matrix (FIM) related to the unknown location parameters according to the received signals, denoted by $\bm{J}_{NF}$. The CRB, which gives the localization performance limit, is then derived from the FIM.
As proved in [24], FIM is given by[35]
$[\bm{J}_{NF}]_{i,j}=E\left\{-\frac{\partial^{2}\ln f(\bm{y}_{k};\bm{p}_{k},\bm{Q})}{\partial p_{i}\partial p_{j}}\right\},$ (59)
where $f(\bm{y}_{k};\bm{p}_{k},\bm{Q})$ is the likelihood function of the random vector $\bm{y}_{k}$ conditioned on $\bm{p}_{k}$ and $\bm{Q}$. Here we assume the transmitted signals are deterministic. Therefore the $i$-th received signal $y^{(i)}_{k}$ can be assumed to be independent of each other and follow complex Gaussian distribution, i.e., $f(y^{(i)}_{k};\bm{p}_{k},\bm{Q})\sim\mathcal{CN}(\mu^{(i)},\sigma^{2})$, where $\mu^{(i)}=(\bm{\beta}^{(i)})^{T}\bm{h}_{k}s$ is the noise-free received signal. Therefore, we have $f(\bm{y}_{k})=\prod_{i=1}^{c+1}f(y^{(i)}_{k})$. Then, in the NF region, the FIM $\bm{J}_{NF}$ can be rewritten as
$[\bm{J}_{NF}]_{i,j}=\dfrac{2}{\sigma^{2}}\sum_{m=1}^{c+1}Re\left\{\dfrac{\partial(\mu^{(m)})^{H}}{\partial p_{i}}\dfrac{\partial\mu^{(m)}}{\partial p_{j}}\right\},$ (60)
where $\frac{\partial\mu^{(m)}}{\partial p_{i}}=(\bm{\beta}^{(m)})^{T}\frac{\partial\bm{h}^{NF}_{k}}{\partial p_{i}}s$. The $n$-th entry of $\frac{\partial\bm{h}^{NF}_{k}}{\partial p_{i}}$ can be expressed as
$\left[\dfrac{\partial\bm{h}^{NF}_{k}}{\partial p_{i}}\right]_{n}=\alpha_{k}\left(-j\dfrac{2\pi}{\lambda}\right)\exp\left(-j\dfrac{2\pi}{\lambda}d_{kn}^{t}\right)\dfrac{\partial d_{kn}^{t}}{\partial p_{i}}.$ (61)
Next, we take the inverse of $\bm{J}_{NF}$ and obtain the CRBs. Hence, the proof of proposition IV-A1 is accomplished.
Appendix B Proof of Proposition IV-A2
Similar to the proof of proposition 2, the FIM $\bm{J}_{FF}$ in the FF region can be written as[35]
$[\bm{J}_{FF}]_{i,j}=\dfrac{2}{\sigma^{2}}\sum_{i=1}^{c+1}Re\left\{\dfrac{\partial(\mu^{(i)})^{H}}{\partial p_{i}}\dfrac{\partial\mu^{(i)}}{\partial p_{j}}\right\},$ (62)
where $\frac{\partial\mu^{(i)}}{\partial p_{i}}=(\bm{\beta}^{(i)})^{T}\frac{\partial\bm{h}^{FF}_{k}}{\partial p_{i}}s$. In the FF region, we have
$\left[\dfrac{\partial\bm{h}^{FF}_{k}}{\partial\theta_{k}}\right]_{n}=$ $h_{k0n}^{t}\left(-j\dfrac{2\pi}{\lambda}\right)(-y_{n}\cos\theta_{k}\sin\phi_{k}+z_{n}\sin\theta_{k}),$ (63)
$\left[\dfrac{\partial\bm{h}^{FF}_{k}}{\partial\phi_{k}}\right]_{n}=$ $h_{k0n}^{t}\left(-j\dfrac{2\pi}{\lambda}\right)(-y_{n}\sin\theta_{k}\cos\phi_{k}),$ (64)
where $h_{k0n}^{t}$ is the $n$-th element in the direct user-RIS path.
Next, we take the inverse of $\bm{J}_{FF}$ and obtain the CRBs. Hence, the proof of proposition IV-A2 is accomplished.
Appendix C The Differentials of CRB
We denote the element in the $i$-th row, $j$-th column of $\bm{J}$ as $J_{ij}$. In order to find the differentials of CRBs, we need to find the differentials of $J_{ij}$. For simplicity, here we use $\bm{\beta}$ to substitute $\bm{\beta}^{(k+1)}$. $J_{ij}$ can be considered as a function of two independent complexed-valued variables $\bm{\beta}$ and $\bm{\beta^{\ast}}$. The complex differentials of $J_{ij}$ can be expressed as
$dJ_{ij}=\left(\dfrac{\partial J_{ij}}{\partial\bm{\beta}}\right)d\bm{\beta}+\left(\dfrac{\partial J_{ij}}{\partial\bm{\beta^{\ast}}}\right)d\bm{\beta}^{\ast}.$ (65)
The complex differentials of $J_{ij}$ are given as
$dJ_{ij}=$ $\dfrac{1}{\sigma^{2}}\left[d\left(s^{H}\left(\dfrac{\partial\bm{h}}{\partial p_{i}}\right)^{H}\bm{\beta}^{\ast}\bm{\beta}^{T}\dfrac{\partial\bm{h}}{\partial p_{j}}s\right)\right.$
$+\left.d\left(s^{T}\left(\dfrac{\partial\bm{h}}{\partial p_{i}}\right)^{T}\bm{\beta}\bm{\beta}^{H}\left(\dfrac{\partial\bm{h}}{\partial p_{j}}\right)^{\ast}s^{\ast}\right)\right].$ (66)
By substitude $\frac{\partial\bm{\beta}^{\ast}\bm{\beta}^{T}}{\partial\bm{\beta}_{n}^{\ast}}$ and $\frac{\partial\bm{\beta}\bm{\beta}^{H}}{\partial\bm{\beta}_{n}^{\ast}}$ into (C), we can obtain
$\dfrac{\partial J_{ij}}{\partial\beta_{n}^{\ast}}=\dfrac{|s|^{2}}{\sigma^{2}}\left[\left(\dfrac{\partial h_{n}}{\partial p_{i}}\right)^{\ast}\sum_{m=1}^{N}\dfrac{\partial h_{m}}{\partial p_{j}}\beta_{m}+\right.$
$\left.\left(\dfrac{\partial h_{n}}{\partial p_{j}}\right)^{\ast}\sum_{m=1}^{N}\dfrac{\partial h_{m}}{\partial p_{i}}\beta_{m}\right].$ (67)
The derivatives of CRBs with respect to $\bm{\beta}^{\ast}$ are given by
$\dfrac{\partial D_{i}}{\partial\bm{\beta}^{*}}$ $=\left(\dfrac{\partial\bm{J}^{-1}}{\partial\bm{\beta}^{*}}\right)_{i,i}=\left(-\bm{J}^{-1}\dfrac{\partial\bm{J}}{\partial\bm{\beta}^{*}}\bm{J}^{-1}\right)_{i,i}.$ (68)